Find the equation of the line containing (4,-3)that is parallel to the line 6x-4y=1
Found 2 solutions by nerdybill, Emsley:
Answer by nerdybill(7384)

Find the equation of the line containing (4,-3)that is parallel to the line 6x-4y=1
.
slope of:
6x-4y=1
-4y=-6x+1
4y=6x-1
y=(6/4)x-1/4
y=(3/2)x-1/4
slope is 3/2
a parallel line must have the same slope.
.
Now, we have the slope (3/2) and one point (4,-3)
plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (3/2)(x - 4)
y + 3 = (3/2)x - (3/2)4
y + 3 = (3/2)x - (3)2
y + 3 = (3/2)x - 6
y = (3/2)x - 9






Answer by Emsley(35)

If it is parallel to 6x-4y=1 its equation will be 6x-4y=k (so that it has same gradient). (4,-3) is on this line so 24+12=k So k=36
So equation is 6x-4y=36